OpenStudy (anonymous):

find exact solution 100 to the x-1 =1000 to the 2x-1

OpenStudy (anonymous):

do you mean 100x-1=1000? Then 2x-1 would equal 20 and x =11

OpenStudy (anonymous):

\[100^{x-1} = 1000^{2x-1}\]

OpenStudy (anonymous):

x=1/4

OpenStudy (anonymous):

how did you figure it out sjano

OpenStudy (anonymous):

Drew it as a graph. :P

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

100^x−1 as one line 1000^2x-1 as line number 2. And they hit eachother at x=0.25 which is 1/4

OpenStudy (anonymous):

you have me really confused sorry

OpenStudy (anonymous):

 Kinda like that